# Recurrent Memory for Online Interdomain Gaussian Processes

### v2

@@ Abstract
We propose a novel online Gaussian process (GP) model that is capable of capturing long-term memory in sequential data in an online regression setting. Our model, Online HiPPO Sparse Variational Gaussian Process Regression (OHSGPR), leverages the HiPPO (High-order Polynomial Projection Operators) framework, which is popularized in the RNN domain due to its long-range memory modeling capabilities. We interpret the HiPPO time-varying orthogonal projections as inducing variables with time-dependent orthogonal polynomial basis functions, which allows the SGPR inducing points to memorize the process history. We show that the HiPPO framework fits naturally into the interdomain GP framework and demonstrate that the kernel matrices can also be updated online in a recurrence form based on the ODE evolution of HiPPO. We evaluate our method on time series regression tasks, showing that it outperforms the existing online GP method in terms of predictive performance and computational efficiency.
@@ 1 Introduction
Gaussian processes (GPs) are a popular choice for modeling time series due to their functional expressiveness and uncertainty quantification abilities (Roberts et al., 2013; Fortuin et al., 2020). However, GPs are computationally expensive and memory intensive, with cubic and quadratic complexities, respectively. In online regression settings, such as weather modeling, the number of time steps can be very large, quickly making GPs infeasible. Although variational approximations, such as utilizing sparse inducing points (SGPR (Titsias, 2009); SVGP (Hensman et al., 2013; 2015a)) and Markovian GPs (Wilkinson et al., 2021), have been proposed to address the computational complexity, it would still be prohibitive to re-fit the GP model from scratch every time new data arrives.
Bui et al. (2017) proposes an online sparse GP (OSGPR) learning method that sequentially updates the GP posterior distribution only based on the newly arrived data. However, as indicated in their paper, their models may not maintain the memory of the previous data, as the inducing points will inevitably shift as new data arrive. This is a major drawback, as their models may not model long-term memory unless using a growing number of inducing points.
@@
In deep learning, as an alternative to Transformers (Vaswani, 2017), significant works on state space models (SSMs) have been proposed to model long-term memory in sequential data. Originally proposed to instill long-term memory in recurrent neural networks, the HiPPO (High-order Polynomial Projection Operators) framework (Gu et al., 2020) provides mathematical foundations for compressing continuous-time signals into memory states through orthogonal polynomial projections. HiPPO is the basis for the state-of-the-art SSMs, such as the structured state space sequential (S4) model (Gu et al., 2022) and Mamba (Gu & Dao, 2023; Dao & Gu, 2024). HiPPO is computationally efficient and exhibits strong performance in long-range memory tasks.
@@
Inspired by HiPPO, we propose Online HiPPO SGPR (OHSGPR), by applying the HiPPO framework to SGPR in order to leverage the long-range memory modeling capabilities. Our method interprets the HiPPO time-varying orthogonal projections as inducing variables of an interdomain SGPR (Lázaro-Gredilla & Figueiras-Vidal, 2009; Leibfried et al., 2020; Van der Wilk et al., 2020), where the basis functions are time-dependent orthogonal polynomials. We show that we are able to significantly resolve the memory-loss issue in OSGPR, thereby opening up the possibility of applying GPs to long-term time series regression tasks. In more detail, we make the following contributions:
@@
  • We demonstrate that HiPPO integrates into the interdomain GPs by interpreting the HiPPO time-varying orthogonal projections as inducing variables with time-dependent orthogonal polynomial basis functions. This allows the inducing variables to memorize the process history, capturing long-term memory in the data.
@@
  • We show that the kernel matrices can leverage the efficient ODE evolution of the HiPPO framework, bringing an extra layer of computational efficiency to OHSGPR.
  • We demonstrate OHSGPR on time series regression tasks, showing that it outperforms vanilla OSGPR in terms of predictive performance and computational efficiency.
@@ 2 Background
In this section, we provide a brief overview of GPs, inducing point methods, and online learning with GPs. In addition, we review the HiPPO method, which is the basis of our proposed method.
@@ 2.1 Gaussian processes
Let $\mathcal{X}$ be the input space. For time series data, $\mathcal{X}=[0,\infty)$, the set of non-negative real numbers. A Gaussian process (GP) $f\sim\mathcal{GP}(0,k)$ is defined with a covariance function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. It has the property that for any finite set of input points $\mathbf{X}=[x_{1},\ldots,x_{n}]^{\intercal}$, the random vector $\mathbf{f}\equiv f(\mathbf{X})=[f(x_{1}),\ldots,f(x_{n})]^{\intercal}\sim\mathcal{N}(0,\mathbf{K}_{\mathbf{f}\mathbf{f}})$, where $\mathbf{K}_{\mathbf{f}\mathbf{f}}$ is the kernel matrix with entries $[k(\mathbf{X},\mathbf{X})]_{ij}\equiv[\mathbf{K}_{\mathbf{f}\mathbf{f}}]_{ij}=k(x_{i},x_{j})$. For notational convenience and different sets of input points $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, we denote the kernel matrix as $\mathbf{K}_{\mathbf{f}_{1}\mathbf{f}_{2}}$ or $k(\mathbf{X}_{1},\mathbf{X}_{2})$. The computational and memory complexities of obtaining the GP posterior on $\mathbf{X}_{1}$ scale cubically and quadratically respectively, according to $n_{1}=|\mathbf{X}_{1}|$. Given responses $\mathbf{y}$ and inputs $\mathbf{X}$, a probabilistic model can be defined as $y_{i}\sim p(y_{i}\mid f(x_{i}))$ with a GP prior $f\sim\mathcal{GP}(0,k)$, where $p(y_{i}\mid f(x_{i}))$ is the likelihood distribution. However, for non-conjugate likelihoods, the posterior distribution is intractable, and approximate inference methods are required, such as, but not limited to, variational inference (Titsias, 2009; Hensman et al., 2013; 2015a) and Markov chain Monte Carlo (MCMC) (Hensman et al., 2015b).
@@ 2.2 Variational inference and interdomain Gaussian processes
To address the intractability and cubic complexity of GPs, Sparse Variational Gaussian Processes (SGPR (Titsias, 2009); SVGP (Hensman et al., 2013; 2015b)) cast the problem as an optimization problem. By introducing $M$ inducing points $\mathbf{Z}\in\mathcal{X}^{M}$ that correspond to $M$ inducing variables $\mathbf{u}=[f(\mathbf{z}_{1}),\ldots,f(\mathbf{z}_{M})]^{\intercal}$, the variational distribution $q(\mathbf{f},\mathbf{u})$ is defined as $q(\mathbf{f},\mathbf{u}):=p(\mathbf{f}\mid\mathbf{u})q_{\theta}(\mathbf{u})$, where $q_{\theta}(\mathbf{u})$ is the variational distribution of the inducing variables with parameters $\theta$. Then, the evidence lower bound (ELBO) is defined as
@@
When the likelihood is conjugate Gaussian, the ELBO can be optimized in closed form and $\mathbf{m}_{\mathbf{u}}$ and $\mathbf{S}_{\mathbf{u}}$ can be obtained in closed form (SGPR; Titsias (2009)). In addition to setting the inducing variables as the function values, interdomain GPs (Lázaro-Gredilla & Figueiras-Vidal, 2009) propose to generalize the inducing variables to $u_{m}:=\int f(x)\phi_{m}(x)\mathrm{d}x$, where $\phi_{m}(x)$ are basis functions, to allow for further flexibility. This yields $[\mathbf{K}_{f\mathbf{u}}]_{m}=\int k(x,x^{\prime})\phi_{m}(x^{\prime})\mathrm{d}x^{\prime}$ and $[\mathbf{K}_{\mathbf{u}\mathbf{u}}]_{nm}=\iint k(x,x^{\prime})\phi_{n}(x)\phi_{m}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime}$.
We see that the interdomain SGPR bypasses the selection of the inducing points $\mathbf{Z}\in\mathbb{R}^{M}$, and reformulates it with the selection of the basis functions $\phi_{i}$. The basis functions dictate the structure of the kernel matrices, which in turn modulate the function space of the GP approximation. In contrast, SGPR relies on the inducing points $\mathbf{Z}$, which can shift locations according to the training data. Some examples of basis functions include Fourier basis functions (Hensman et al., 2018) and the Dirac delta function $\delta_{\mathbf{z}_{m}}$, the latter recovering the standard SGPR inducing variables.
@@ 2.3 Online Gaussian processes
In this paper, we focus on online learning with GPs, where data arrives sequentially in batches $(\mathbf{X}_{t_{1}},\mathbf{y}_{t_{1}}),(\mathbf{X}_{t_{2}},\mathbf{y}_{t_{2}}),\ldots$ etc. In the time series regression setting, the data arrives in intervals of $(0,t_{1}),(t_{1},t_{2}),\ldots$ etc. The online GP learning problem is to sequentially update the GP posterior distribution as data arrives. Suppose that we have already obtained $p_{t_{1}}(y|f)p_{t_{1}}(f|\mathbf{u}_{t_{1}})q_{t_{1}}(\mathbf{u}_{t_{1}})$ of the likelihood and variational approximation (with inducing points $\mathbf{Z}_{t_{1}}$), from the first data batch $(\mathbf{X}_{t_{1}},\mathbf{y}_{t_{1}})$. Online SGPR (OSGPR; Bui et al. (2017)) utilizes the online learning ELBO
$\sum_{i=1}^{n_{t_{2}}}\mathop{\mathbb{E}}_{q_{t_{2}}(f_{i})}\left[\log p_{t_{2}}(y_{i}\mid f_{i})\right]+\text{KL}\left(\tilde{q}_{t_{2}}(\mathbf{u}_{t_{1}})\,\middle\|\,p_{1}(\mathbf{u}_{t_{1}})\right)$ (3)
$\quad-\text{KL}\left(\tilde{q}_{t_{2}}(\mathbf{u}_{t_{2}})\,\middle\|\,q_{t_{1}}(\mathbf{u}_{t_{1}})\right)-\text{KL}\left(q_{t_{2}}(\mathbf{u}_{t_{2}})\,\middle\|\,p_{t_{2}}(\mathbf{u}_{t_{2}})\right),$
where $y_{i}\in\mathbf{y}_{t_{2}}$ for $i=1,\ldots,n_{t_{2}}$ and $\tilde{q}_{t_{2}}(\mathbf{u}_{t_{1}}):=\int p_{t_{2}}(\mathbf{u}_{t_{1}}|\mathbf{u}_{t_{2}})q_{t_{2}}(\mathbf{u}_{t_{2}})\mathrm{d}\mathbf{u}_{t_{2}}$. Note that subscripts $()_{t_{1}}$ are utilizing hyperparameters from the first batch, and $()_{t_{2}}$ are hyperparameters from the second batch that are to be optimized with the ELBO. Theoretically, Bui et al. (2017) show that the OSGPR method achieves the same posterior distribution as the batch SGPR method when the inducing points are fixed. Unfortunately, for online time series regression, OSGPR may not capture the long-term memory in the time series data since as new data arrives, it is not guaranteed that the inducing points after optimization can sufficiently cover all the previous tasks’ data domains.
@@ 2.4 HiPPO: recurrent memory with optimal polynomial projections
The HiPPO framework (Gu et al., 2020) provides mathematical foundations for compressing continuous-time signals into finite-dimensional memory states through optimal polynomial projections. Given a time series $y(t)$, HiPPO maintains a memory state $c(t)\in\mathbb{R}^{M}$ that optimally approximates the historical signal $\{y(s)\}_{s\leq t}$. The framework consists of a time-dependent measure $\omega^{(t)}(s)$ over $(-\infty,t]$ that defines input importance, along with normalized polynomial basis functions $\{g_{n}^{(t)}(s)\}_{n=0}^{M-1}$ that are orthonormal under $\omega^{(t)}(s)$, satisfying $\int_{-\infty}^{t}g_{m}^{(t)}(s)g_{n}^{(t)}(s)\omega^{(t)}(s)\mathrm{d}s=\delta_{mn}$. The historical signal is encoded through projection coefficients given by
$c_{n}(t)=\int_{-\infty}^{t}y(s)g_{n}^{(t)}(s)\omega^{(t)}(s)\mathrm{d}s.$ (4)
This yields the approximation $y(s)\approx\sum_{n=0}^{M-1}c_{n}(t)g_{n}(s)$ for $s\in(-\infty,t]$, minimizing the $L^{2}$-error $\int_{-\infty}^{t}\|y(s)-\sum_{n}c_{n}(t)g_{n}(s)\|^{2}\omega^{(t)}(s)\mathrm{d}s$.
Differentiating $\mathbf{c}(t):=[c_{0}(t),\ldots,c_{M-1}(t)]^{\intercal}$ induces a linear ordinary differential equation $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{c}(t)=\mathbf{A}(t)\mathbf{c}(t)+\mathbf{B}(t)y(t)$ with matrices $\mathbf{A}(t),\mathbf{B}(t)$ encoding measure-basis dynamics. Discretization yields the recurrence $\mathbf{c}_{t}=\mathbf{A}_{t}\mathbf{c}_{t-1}+\mathbf{B}_{t}y_{t}$ enabling online updates. The structured state space sequential (S4) model (Gu et al., 2022) extends HiPPO with trainable parameters and convolutional kernels, while Mamba (Gu & Dao, 2023; Dao & Gu, 2024) introduces hardware-aware selective state mechanisms, both leveraging HiPPO for efficient long-range memory modeling.
HiPPO supports various measure-basis configurations (Gu et al., 2020; 2023). A canonical instantiation, HiPPO-LegS, uses a uniform measure $\omega^{(t)}(s)=\frac{1}{t}\mathbf{1}_{[0,t]}(s)$ with scaled Legendre polynomials adapted to $[0,t]$, $g_{n}^{(t)}(s)=(2m+1)^{1/2}P_{m}\left(\frac{2s}{t}-1\right)$. This uniform measure encourages HiPPO-LegS to keep the whole past in memory.

### v4

@@ Abstract
We propose a novel online Gaussian process (GP) model that is capable of capturing long-term memory in sequential data in an online learning setting. Our model, Online HiPPO Sparse Variational Gaussian Process (OHSVGP), leverages the HiPPO (High-order Polynomial Projection Operators) framework, which is popularized in the RNN domain due to its long-range memory modeling capabilities. We interpret the HiPPO time-varying orthogonal projections as inducing variables with time-dependent orthogonal polynomial basis functions, which allows the SVGP inducing variables to memorize the process history. We show that the HiPPO framework fits naturally into the interdomain GP framework and demonstrate that the kernel matrices can also be updated online in a recurrence form based on the ODE evolution of HiPPO. We evaluate OHSVGP with online prediction for 1D time series, continual learning in discriminative GP model for data with multidimensional inputs, and deep generative modeling with sparse Gaussian process variational autoencoder, showing that it outperforms existing online GP methods in terms of predictive performance, long-term memory preservation, and computational efficiency.
footnotetext: *Equal contribution.footnotetext: Source Code: https://github.com/harrisonzhu508/HIPPOSVGP.
@@ 1 Introduction
Gaussian processes (GPs) are popular choices for modeling time series due to their functional expressiveness and uncertainty quantification abilities (Roberts et al., 2013; Fortuin et al., 2020). However, GPs are computationally expensive and memory intensive, with cubic and quadratic complexities, respectively. In online regression settings, such as weather modeling, the number of time steps can be very large, quickly making GPs infeasible. Although variational approximations, such as utilizing sparse inducing points (SGPR (Titsias, 2009); SVGP (Hensman et al., 2013, 2015a)) and Markovian GPs (Särkkä and Solin, 2019; Wilkinson et al., 2021), have been proposed to address the computational complexity, it would still be prohibitive to re-fit the GP model from scratch every time new data arrives. Bui et al. (2017) proposed an online sparse variational GP (OSVGP) learning method that sequentially updates the GP posterior distribution only based on the newly arrived data. However, as indicated in their paper, their models may not maintain the memory of the previous data, as the inducing points will inevitably shift as new data arrive. This is a major drawback, as their models may not model long-term memory unless using a growing number of inducing points.
@@
In deep learning, as an alternative to Transformers (Vaswani et al., 2017), significant works on state space models (SSMs) have been proposed to model long-term memory in sequential data. Originally proposed to instill long-term memory in recurrent neural networks, the HiPPO (High-order Polynomial Projection Operators) framework (Gu et al., 2020) provides mathematical foundations for compressing continuous-time signals into memory states through orthogonal polynomial projections. HiPPO is computationally efficient and exhibits strong performance in long-range memory tasks, and forms the basis for the state-of-the-art SSMs, e.g., structured state space sequential (S4) model (Gu et al., 2022) and Mamba (Gu and Dao, 2023; Dao and Gu, 2024).
@@
Inspired by HiPPO, we propose Online HiPPO SVGP (OHSVGP), by applying the HiPPO framework to SVGP in order to leverage the long-range memory modeling capabilities. Our method interprets the HiPPO time-varying orthogonal projections as inducing variables of an interdomain SVGP (Lázaro-Gredilla and Figueiras-Vidal, 2009; Leibfried et al., 2020; Van der Wilk et al., 2020), where the basis functions are time-dependent orthogonal polynomials. We show that we are able to significantly resolve the memory-loss issue in OSVGP, thereby opening up the possibility of applying GPs to long-term online learning tasks. In summary, our contributions include:
@@
  • (Section 3) We demonstrate that HiPPO integrates into the interdomain GPs by interpreting the HiPPO projections as inducing variables with time-dependent orthogonal polynomial basis functions. This allows the inducing variables to compress historical data, capturing long-term information.
@@
  • (Section 3.25.1) We show that the kernel matrices can leverage the efficient ODE evolution of the HiPPO framework, bringing an extra layer of computational efficiency to OHSVGP.
  • (Section 5) We demonstrate OHSVGP on a variety of online/continual learning tasks including time series prediction, continual learning on UCI benchmarks, and continual learning in Gaussian process variational autoencoder, showing that it outperforms other online sparse GP baselines in terms of predictive performance, long-term memory preservation, and computational efficiency.
@@ 2 Background
In this section, we provide a brief overview of GPs, inducing point methods, online learning with GPs, and Gaussian process variational autoencoders. In addition, we review the HiPPO method, which is the basis of our proposed method.
@@ 2.1 Gaussian processes
Let $\mathcal{X}$ be the input space. For time series data, $\mathcal{X}=[0,\infty)$, the set of non-negative real numbers. A Gaussian process (GP) $f\sim\mathcal{GP}(0,k)$ is defined with a covariance function $k:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. It has the property that for any finite set of input points $\mathbf{X}=[x_{1},\ldots,x_{n}]^{\intercal}$, the random vector $\mathbf{f}\equiv f(\mathbf{X})=[f(x_{1}),\ldots,f(x_{n})]^{\intercal}\sim\mathcal{N}(0,\mathbf{K}_{\mathbf{f}\mathbf{f}})$, where $\mathbf{K}_{\mathbf{f}\mathbf{f}}$ is the kernel matrix with entries $k(\mathbf{X},\mathbf{X})]_{ij}\equiv[\mathbf{K}_{\mathbf{f}\mathbf{f}}]_{ij}=k(x_{i},x_{j})$. For notational convenience and different sets of input points $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$, we denote the kernel matrix as $\mathbf{K}_{\mathbf{f}_{1}\mathbf{f}_{2}}$ or $k(\mathbf{X}_{1},\mathbf{X}_{2})$. The computational and memory complexities of obtaining the GP posterior on $\mathbf{X}_{1}$ scale cubically and quadratically respectively, according to $n_{1}=|\mathbf{X}_{1}|$. Given responses $\mathbf{y}$ and inputs $\mathbf{X}$, a probabilistic model can be defined as $y_{i}\sim p(y_{i}\mid f(x_{i}))$ with a GP prior $f\sim\mathcal{GP}(0,k)$, where $p(y_{i}\mid f(x_{i}))$ is the likelihood distribution. However, for non-conjugate likelihoods, the posterior distribution is intractable, and approximate inference methods are required, such as, but not limited to, variational inference (Titsias, 2009; Hensman et al., 2013, 2015a) and Markov chain Monte Carlo (MCMC) (Hensman et al., 2015b).
@@ 2.2 Variational inference and interdomain Gaussian processes
To address the intractability and cubic complexity of GPs, Sparse Variational Gaussian Processes (SVGP; (Titsias, 2009; Hensman et al., 2013, 2015b)) cast the problem as an optimization problem. By introducing $M$ inducing points $\mathbf{Z}\in\mathcal{X}^{M}$ that correspond to $M$ inducing variables $\mathbf{u}=[f(\mathbf{z}_{1}),\ldots,f(\mathbf{z}_{M})]^{\intercal}$, the variational distribution $q(\mathbf{f},\mathbf{u})$ is defined as $q(\mathbf{f},\mathbf{u}):=p(\mathbf{f}\mid\mathbf{u})q_{\theta}(\mathbf{u})$, where $q_{\theta}(\mathbf{u})$ is the variational distribution of the inducing variables with parameters $\theta$. Then, the evidence lower bound (ELBO) is defined as
@@
When the likelihood is conjugate Gaussian, the ELBO can be optimized in closed form and $\mathbf{m}_{\mathbf{u}}$ and $\mathbf{S}_{\mathbf{u}}$ can be obtained in closed form (SGPR; Titsias (2009)). In addition to setting the inducing variables as the function values, interdomain GPs (Lázaro-Gredilla and Figueiras-Vidal, 2009) propose to generalize the inducing variables to $u_{m}:=\int f(x)\phi_{m}(x)\mathrm{d}x$, where $\phi_{m}(x)$ are basis functions, to allow for further flexibility. This yields $\mathbf{K}_{f\mathbf{u}}]_{m}=\int k(x,x^{\prime})\phi_{m}(x^{\prime})\mathrm{d}x^{\prime}$ and $\mathbf{K}_{\mathbf{u}\mathbf{u}}]_{nm}=\iint k(x,x^{\prime})\phi_{n}(x)\phi_{m}(x^{\prime})\mathrm{d}x\mathrm{d}x^{\prime}$.
We see that the interdomain SVGP bypasses the selection of the inducing points $\mathbf{Z}\in\mathbb{R}^{M}$, and reformulates it with the selection of the basis functions $\phi_{i}$. The basis functions dictate the structure of the kernel matrices, which in turn modulate the function space of the GP approximation. In contrast, SVGP relies on the inducing points $\mathbf{Z}$, which can shift locations according to the training data. Some examples of basis functions include Fourier basis functions (Hensman et al., 2018) and the Dirac delta function $\delta_{\mathbf{z}_{m}}$, the latter recovering the standard SVGP inducing variables.
@@ 2.3 Online Gaussian processes
In this paper, we focus on online learning with GPs, where data arrives sequentially in batches $(\mathbf{X}_{t_{1}},\mathbf{y}_{t_{1}}),(\mathbf{X}_{t_{2}},\mathbf{y}_{t_{2}}),\ldots$ etc. For example, in the time series prediction setting, the data arrives in intervals of $(0,t_{1}),(t_{1},t_{2}),\ldots$ etc. The online GP learning problem is to sequentially update the GP posterior distribution as data arrives. Suppose that we have already obtained $p_{t_{1}}(y|f)p_{t_{1}}(f|\mathbf{u}_{t_{1}})q_{t_{1}}(\mathbf{u}_{t_{1}})$ of the likelihood and variational approximation (with inducing points $\mathbf{Z}_{t_{1}}$), from the first data batch $(\mathbf{X}_{t_{1}},\mathbf{y}_{t_{1}})$. Online SVGP (OSVGP; (Bui et al., 2017)) utilizes the online learning ELBO
$\sum_{i=1}^{n_{t_{2}}}\mathop{\mathbb{E}}_{q_{t_{2}}(f_{i})}\left[\log p_{t_{2}}(y_{i}\mid f_{i})\right]+\text{KL}\left(\tilde{q}_{t_{2}}(\mathbf{u}_{t_{1}})\,\middle\|\,p_{t_{1}}(\mathbf{u}_{t_{1}})\right)$ (3)
$\quad-\text{KL}\left(\tilde{q}_{t_{2}}(\mathbf{u}_{t_{1}})\,\middle\|\,q_{t_{1}}(\mathbf{u}_{t_{1}})\right)-\text{KL}\left(q_{t_{2}}(\mathbf{u}_{t_{2}})\,\middle\|\,p_{t_{2}}(\mathbf{u}_{t_{2}})\right),$
where $y_{i}\in\mathbf{y}_{t_{2}}$ for $i=1,\ldots,n_{t_{2}}$ and $\tilde{q}_{t_{2}}(\mathbf{u}_{t_{1}}):=\int p_{t_{2}}(\mathbf{u}_{t_{1}}|\mathbf{u}_{t_{2}})q_{t_{2}}(\mathbf{u}_{t_{2}})\mathrm{d}\mathbf{u}_{t_{2}}$. Unfortunately, with more and more tasks, OSVGP may not capture the long-term memory in the data since as new data arrives, it is not guaranteed that the inducing points after optimization can sufficiently cover all the previous tasks’ data domains.
2.4 Gaussian process variational autoencoders
Gaussian processes can be embedded within a variational autoencoder (VAE; (Kingma and Welling, 2014)) framework, giving rise to the Gaussian process variational autoencoder (GPVAE; (Casale et al., 2018; Fortuin et al., 2020; Ashman et al., 2020; Jazbec et al., 2021; Zhu et al., 2023)). For sparse GPs with inducing variables, Jazbec et al. (2021) introduced the SVGPVAE, which combines the sparse variational GP (SVGP) with the VAE formulation. The likelihood $p(y\mid\varphi_{\theta}(f))$ is parameterized by a decoder network $\varphi_{\theta}$, which takes GP latent draws $f$ as input, together with the variational inducing posterior $q_{\theta}(\mathbf{u}\mid y)$. This posterior, $q_{\theta}(\mathbf{u}\mid\phi(y))$, is parameterized by the encoder network $\phi$. Finally, the latent GP $f$ is typically modeled as a multi-output GP with independent components. GPVAEs have been shown to successfully model high-dimensional time series such as weather data and videos (Zhu et al., 2023; Fortuin et al., 2020). In this work, we consider the SVGPVAE model defined in Jazbec et al. (2021) for one set of our experiments, and the detailed specification of the model and training objective can be found in Appendix D.2.
@@ 2.5 HiPPO: recurrent memory with optimal polynomial projections
The HiPPO framework (Gu et al., 2020) provides mathematical foundations for compressing continuous-time signals into finite-dimensional memory states through optimal polynomial projections. Given a time series $y(t)$, HiPPO maintains a memory state $c(t)\in\mathbb{R}^{M}$ that optimally approximates the historical signal $\{y(x)\}_{x\leq t}$. The framework consists of a time-dependent measure $\omega^{(t)}(x)$ over $(-\infty,t]$ that defines input importance, along with normalized polynomial basis functions $\{g_{n}^{(t)}(x)\}_{n=0}^{M-1}$ that are orthonormal under $\omega^{(t)}(x)$, satisfying $\int_{-\infty}^{t}g_{m}^{(t)}(x)g_{n}^{(t)}(x)\omega^{(t)}(x)\mathrm{d}x=\delta_{mn}$. The historical signal is encoded through projection coefficients given by $c_{n}(t)=\int_{-\infty}^{t}y(x)g_{n}^{(t)}(x)\omega^{(t)}(x)\mathrm{d}x$. This yields the approximation $y(x)\approx\sum_{n=0}^{M-1}c_{n}(t)g_{n}(x)$ for $x\in(-\infty,t]$, minimizing the $L^{2}$-error $\int_{-\infty}^{t}\|y(x)-\sum_{n}c_{n}(t)g_{n}(x)\|^{2}\omega^{(t)}(x)\mathrm{d}x$. Differentiating $\mathbf{c}(t):=[c_{0}(t),\ldots,c_{M-1}(t)]^{\intercal}$ induces a linear ordinary differential equation $\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{c}(t)=\mathbf{A}(t)\mathbf{c}(t)+\mathbf{B}(t)y(t)$ with matrices $\mathbf{A}(t),\mathbf{B}(t)$ encoding measure-basis dynamics. Discretization yields the recurrence of the form $\mathbf{c}_{t}=\mathbf{A}_{t}\mathbf{c}_{t-1}+\mathbf{B}_{t}y_{t}$ enabling online updates. The structured state space sequential (S4) model (Gu et al., 2022) extends HiPPO with trainable parameters and convolutional kernels, while Mamba (Gu and Dao, 2023; Dao and Gu, 2024) introduces hardware-aware selective state mechanisms, both leveraging HiPPO for efficient long-range memory modeling. HiPPO supports various measure-basis configurations (Gu et al., 2020, 2023). A canonical instantiation, HiPPO-LegS, uses a uniform measure $\omega^{(t)}(x)=\frac{1}{t}\mathbf{1}_{[0,t]}(x)$ with scaled Legendre polynomials adapted to $[0,t]$, $g_{n}^{(t)}(x)=(2m+1)^{1/2}P_{m}\left(\frac{2x}{t}-1\right)$. This uniform measure encourages HiPPO-LegS to keep the whole past in memory.